On a frictionless horizontal air table, puck A (mass 0.246 kg) is moving to the right toward puck B (mass 0.667 kg), which is initially at rest. Puck A collides inelastically with puck B. After the collision, puck A has velocity 0.207 m/s to the left, and puck B has velocity 0.532 m/s to the right.

  1. What was the speed of puck A before the collision?
  2. Calculate the initial kinetic energy of puck A.
  3. Calculate the change in total kinetic energy of the system that occurs during the collision.
1
Expert's answer
2021-10-25T10:03:08-0400

1) Use momentum conservation:


"mv=-mv_m+Mv_M,\\
v=1.235\text{ m/s}."

2) The initial energy:


"Ke=0.5mv^2=0.188\text{ J}."

3) The change is the difference between the initial energy and final energies of both pucks:


"∆=Ke-Ke_m-Ke_M=0.0883\text{ J}."